find f of g if f(x)= x-x² and g(x)= x-2
Answer by stanbon(75887)

find f of g if f(x)= x-x² and g(x)= x-2
----
fog(x) = f[x-2] = (x-2)-(x-2)^2
---
x-2 -[x^2-4x+4]
--------
x-2-x^2+4x-4
---
= -x^2 + 5x -6
==================================
Cheers,
Stan H.
================